# Scenario-Based Curriculum Generation for Multi-Agent Autonomous Driving

### v1

@@ Abstract
The automated generation of diverse and complex training scenarios has been an important ingredient in many complex learning tasks. Especially in real-world application domains, such as autonomous driving, auto-curriculum generation is considered vital for obtaining robust and general policies. However, crafting traffic scenarios with multiple, heterogeneous agents is typically considered as a tedious and time-consuming task, especially in more complex simulation environments. In our work, we introduce MATS-Gym, a Multi-Agent Traffic Scenario framework to train agents in CARLA, a high-fidelity driving simulator. MATS-Gym is a multi-agent training framework for autonomous driving that uses partial scenario specifications to generate traffic scenarios with variable numbers of agents. This paper unifies various existing approaches to traffic scenario description into a single training framework and demonstrates how it can be integrated with techniques from unsupervised environment design to automate the generation of adaptive auto-curricula. The code is available at https://github.com/AutonomousDrivingExaminer/mats-gym.
@@ I Introduction
Autonomous driving (AD) stands as a groundbreaking frontier in transportation technology, offering the promise of safer and more efficient roadways. However, achieving fully autonomous vehicles has been a long-standing goal for researchers across generations. As recent advancements enable the deployment of more vehicles in cities worldwide, extensive testing in a diverse range of scenarios becomes essential. For every kilometer driven by AD systems on real-world roads, they traverse several orders of magnitude more in simulation. Thus, access to high-fidelity simulation environments plays a crucial role in advancing the field of AD. Open-source simulation software like CARLA [1] has empowered researchers and practitioners to evaluate and train AD algorithms in realistic environments. Despite this, creating diverse and lifelike traffic scenarios, incorporating varying numbers of agents, remains a laborious and time-consuming task. To tackle this challenge, scenario specification languages like Scenic [2] and OpenSCENARIO [^1] https://www.asam.net/standards/detail/openscenario/ have been introduced. These scenario specification approaches streamline the process of defining logical traffic scenarios and are well-integrated into simulation environments such as CARLA. However, we observe that (1) infrastructure aiding the training and evaluation of AD stacks is scattered across numerous ad hoc implementations, (2) integration of scenario specifications and simulation environments lacks support for training in multi-agent systems and (3) limited support for approaches targeting open-ended curriculum generation. To address these shortcomings, we propose MATS-Gym, an open-source multi-agent training framework that leverages existing traffic scenario-specification approaches for CARLA. Additionally, we demonstrate how MATS-Gym can be used to implement novel auto-curriculum approaches [3] for generating open-ended traffic scenarios with adaptive difficulty. Hence, the key contributions of MATS-Gym are:
@@ II-A Training Frameworks for AD
We compare our approach to other frameworks commonly utilized for training and evaluating AD agents. We identified several popular frameworks and assess them based on several key aspects crucial for our use-case. These aspects include ease and expressivity of scenario specifications, simulation realism, supported traffic types, and integrated training interfaces. Table I summarizes the comparison.
@@
HighwayEnv [4], BARK [5], SMARTS [8], and the CommonRoads Suite [6, 7] offer alternatives to full-fledged game-engine simulations like CARLA. They focus on modeling vehicle dynamics and motion planning, with CommonRoads providing more sophisticated multi-body dynamics compared to HighwayEnv and BARK. While simpler environments are computationally efficient, they lack complex sensor models and realistic observations. These frameworks mainly involve vehicles and lack pedestrian or object modeling found in urban traffic. BARK and CommonRoads allow scriptable scenarios, while HighwayEnv supports custom scenarios through code re-implementation. SMARTS allows to model urban traffic scenarios with traffic lights and traffic signs. However, since the focus is not on graphical realism, those scenarios can not reach the visual diversity of traffic scenes in CARLA, which is necessary to train and evaluate vision based AD stacks.
@@
MetaDrive [9] is a training framework that builds on the Panda3D engine. It offers a wide range of features such as procedural road generation and incorporating scenarios from traffic datasets. However, scenarios can not be specified by behavioral building blocks or predefined sub-scenarios. Moreover, its engine limits the structural variety of environments that can be achieved in CARLA.
@@
DI Drive Gym [10] and MACAD Gym [11] are both training frameworks designed around CARLA. DI Drive Gym focuses on single-agent autonomous driving systems and interfaces with ScenarioRunner scenarios, lacking support for training multi-agent systems or scenario generation via sampling. MACAD Gym, supports multi-agent training but offers limited scenario generation capabilities, allowing only manual specification of initial conditions.
@@
In this comparison, we focus on relevant aspects for learning-based approaches in realistic urban traffic scenarios. We acknowledge our criteria may not be complete and there might be aspects that could be relevant in other use cases. However, we tried to cover a broad spectrum of frameworks that could be used for training and evaluation.
II-B Curriculum Learning
@@
Generating safety-critical scenarios is an active area of research and has been addressed in numerous works in recent years [12]. However, training agents in complex environments requires presenting levels at the right difficulty, matching their capabilities to generate meaningful learning signals. Curriculum Learning (CL) and Auto CL (ACL) methods have emerged to progressively expose agents to more complex environments. While traditional CL involves manually designing curricula, ACL automates this process by dynamically presenting suitable scenarios based on agent performance. For an overview, we refer to [13].
@@
Previous works also applied ACL to AD scenarios [14, 15]. A novel research direction in ACL is Unsupervised Environment Design (UED), where task creation itself is treated as an optimization problem. This is formalized through Underspecified Partially Observable Markov Decision Processes (UPOMDP), allowing for flexible reparameterization of environments during runtime. A UPOMDP is a tuple
$\mathcal{M}=\langle A,O,\Theta,{S}^{\mathcal{M}},{\mathcal{T}}^{\mathcal{M}},{\mathcal{I}}^{\mathcal{M}},{\mathcal{R}}^{\mathcal{M}},\gamma\rangle,$
@@
which differs from a standard POMDP only in having a set of free parameters $\Theta$, which can be chosen at any time during runtime. Recent works have explored different algorithms for UED. Prioritized Level Replay (PLR) utilizes domain randomization to generate parameters and a level replay buffer to resample scenarios proportional to their regret, a measure of suboptimality [16]. PAIRED [17] optimizes the level generator via reinforcement learning (RL) to maximize regret, while REPAIRED [3] combines PLR with the level generation approach from PAIRED in a Dual Curriculum Design (DCD) framework. Other approaches use evolutionary algorithms to optimize the generator and include a level editor for environment modifications [18, 19].
III MATS-Gym: Architecture and Features
MATS-Gym is a multi-agent training and evaluation framework that allows to generate diverse traffic scenarios in CARLA, a high-fidelity traffic simulator. The framework is designed to reconcile scenario execution engines, such as Scenic [2] and ScenarioRunner [20], and training frameworks in the AD research community. We model the interaction between the environment and the agents as a Partially Observable Stochastic Game (POSG) where all agents take actions simultaneously at each timestep. Thus, a natural choice for the environment interface is the PettingZoo API [21], the multi-agent version of OpenAI Gym [22].
III-A Architecture
Figure 2: MATS-Gym Architecture: We build on ScenarioRunner the scenario execution infrastructure. Blue: our contributions, Green: existing infrastructure from ScenarioRunner.
Figure 3: Action abstractions include low-level control, waypoints for local planning, and macro-actions for behavior planning.
We build MATS-Gym around the core of ScenarioRunner, which allows us to leverage their existing infrastructure to specify scenarios, define event observers and execute scenarios, as depicted in Figure 2. Our framework serves as an execution handler and training interface that extends the capabilities of those scenarios to multi-agent environments. Furthermore, we provide a set of wrappers to enhance observation and action spaces in order to facilitate versatile training environments. Additionally, we also provide a task infrastructure that allows to easily define reward functions and termination conditions. One of the design goals of MATS-Gym is interoperability with current scenario engines for CARLA. Compatibility with ScenarioRunner and thus OpenSCENARIO is inherently built in. As a consequence, our framework can be also used as a training and evaluation framework for the CARLA AD Challenge [^2] https://leaderboard.carla.org/challenge/. Furthermore, we connect another complementary scenario specification engine for CARLA by implementing an adapter for Scenic, a probabilistic programming language for building complex scenarios in various domains. Scenic comes with a rich set of building blocks for specifying distributions over traffic scenarios that adhere to predefined constraints.
III-B Observations, Actions and Tasks
@@
MATS-Gym offers comprehensive infrastructure for agent state retrieval, providing ground truth information such as position, velocity vectors, and key traffic events like violations and collisions. Its modular design allows for flexible specification of additional observation types. It offers a versatile infrastructure for configuring sensor suites for agents, akin to the CARLA AD Challenge, supporting various sensors such as cameras, LiDAR, and Radar. It extends the CARLA BirdEyeView [23] framework for bird’s-eye view observations, constructing occupancy gridmaps around agents encoding lane markings, traffic lights, signs, and other road users. Agents can also access road network information and vector-based representations of the road layout. We account for different type of tasks by providing action spaces at various levels of abstraction. On the lowest level, agents can issue throttle, breaking and steering commands at high frequencies. On a higher level, agents can instead provide a target waypoint which is tracked by a PID controller for smooth driving trajectories. We also provide a discrete action space for high-level driving commands, suitable for route planning, behavior prediction, or multi-agent interaction. Fig. 3 visualizes the available action spaces. Furthermore, MATS-Gym allows for defining custom tasks and offers pre-defined tasks related to autonomous driving, including route following, infraction avoidance, and driving comfort, crucial for designing reward functions in reinforcement learning approaches.
IV Scenario Based Curriculum Generation
@@
In the following, we will outline our approach to use Scenic’s parametrizable scenario specifications as a way to build an adaptive scenario generation procedure. We adopt a dual-curriculum design algorithm based on REPAIRED [3], a recent algorithm combining PLR with an adaptive level generator. At each timestep, we either generate a new level by sampling environment parameters from $\tilde{\pi}$ or we sample a replay level from our PLR buffer. The buffer samples levels with a probability proportional to their regret and a staleness metric, as described in the original approach. Then, we sample a trajectory from the environment to update the student policy $\pi$ and to compute the estimated regret for the current level, which is then used to update the buffer and the environment generator. Algorithm 1 outlines this approach.
Differently than the original, we do not use an antagonist agent to estimate the regret, due to computational constraints. Instead, we use the Maximum Monte Carlo [3] regret estimator, formally described as:
$\frac{1}{T}\sum_{t=0}^{T}R_{\text{max}}-V(s_{t}).$ (1)
This regret formulation compares the maximum achieved return with the average state value for each episode. We choose this alternative to avoid training another policy just for regret estimation, which would entail twice as much simulation steps. Furthermore, for the level generator optimization, we use the tree-structured Parzen estimator algorithm, implemented in the hyperparameter optimization framework Optuna [24] instead of RL.
@@
1:  Input: policy $\pi$, environment generator $\tilde{\pi}$, buffer $\Lambda$
@@
5:        Generate environment parameters $\theta\sim\tilde{\pi}$.
@@
8:        Sample replay level, $\theta\sim\Lambda$.
@@
10:     Collect trajectory $\tau$ using policy $\pi$.
11:     Compute regret $R$ using Eq. 1.
12:     Update policy $\pi$ with $\tau$.
13:     Update level buffer: $\Lambda\leftarrow(\theta,R)$.
14:     if level was generated then
15:        Update environment generator $\tilde{\pi}$ with $R$.

### v2

# Scenario-Based Curriculum Generation for Multi-Agent Driving

@@ Abstract
The automated generation of diversified training scenarios has been an important ingredient in many complex learning tasks, especially in real-world application domains such as autonomous driving, where auto-curriculum generation is considered vital for obtaining robust and general policies. However, crafting traffic scenarios with multiple, heterogeneous agents, is typically considered as a tedious and time-consuming task, especially in more complex simulation environments. To this end, we introduce MATS-Gym, a multi-agent training framework for autonomous driving, that uses partial-scenario specifications to generate traffic scenarios with a variable number of agents which are executed in CARLA, a high-fidelity driving simulator. MATS-Gym reconciles scenario execution engines, such as Scenic and ScenarioRunner, with established multi-agent training frameworks where the interaction between the environment and the agents is modeled as a partially-observable stochastic game. Furthermore, we integrate MATS-Gym with techniques from unsupervised environment design to automate the generation of adaptive auto-curricula, which is the first application of such algorithms to the domain of autonomous driving. The code is available at https://github.com/AutonomousDrivingExaminer/mats-gym.
@@ I Introduction
Autonomous driving (AD) is a groundbreaking frontier of transportation technology, offering the promise of safer, and more efficient roadways. However, full autonomy is still a long-standing, but not-yet achieved goal for researchers across generations. As recent advancements enable the deployment of more vehicles in cities worldwide, extensive testing in a diverse range of scenarios becomes essential. For every kilometer driven by AD vehicles on real-world roads, they traverse several orders of magnitude more in simulation. Thus, access to high-fidelity simulation environments plays a crucial role in advancing the field of AD.
Open-source simulation software like CARLA [1] has already empowered researchers and practitioners to evaluate and train AD algorithms in realistic environments. Despite this, creating diverse and lifelike traffic scenarios incorporating varying numbers of agents remains a laborious and time-consuming task. To tackle this challenge, scenario specification languages like Scenic [2, 3] and OpenSCENARIO [^1] https://www.asam.net/standards/detail/openscenario/ have been introduced. These scenario specification approaches streamline the process of defining logical traffic scenarios and are well-integrated into simulation environments such as CARLA. However, we observe that: (1) infrastructure aiding the training and evaluation of AD stacks is scattered across numerous ad hoc implementations, (2) integration of scenario specifications and simulation environments lacks support for training in multi-agent systems, and (3) approaches targeting open-ended curriculum generation have limited support. To address these shortcomings, we propose MATS-Gym, an open-source multi-agent training framework that leverages existing traffic scenario-specification approaches for CARLA. Additionally, we demonstrate how MATS-Gym can be used to implement novel auto-curriculum approaches [4] for generating open-ended traffic scenarios with adaptive difficulty. The key contributions of MATS-Gym are:
@@
    A versatile and flexible configuration infrastructure for our framework that supports multiple scenario description and sampling solutions, sensor models and levels of abstraction for defining action and state spaces.
  3. 3.
@@ II Related Work
Training Frameworks for AD. We compare our approach to other frameworks commonly utilized for training and evaluating AD agents. We identified several popular frameworks and assess them based on several key aspects crucial for our use-case. These aspects include ease and expressivity of scenario specifications, simulation realism, supported traffic types, and integrated training interfaces. Table I summarizes the comparison.
@@
HighwayEnv [5], BARK [6], SMARTS [9], and the CommonRoads Suite [7, 8] offer alternatives to full-fledged game-engine simulations like CARLA. They focus on modeling vehicle dynamics and motion planning, with CommonRoads providing more sophisticated multi-body dynamics compared to HighwayEnv and BARK. While simpler environments are computationally efficient, they lack complex sensor models and realistic observations. These frameworks mainly involve vehicles and lack pedestrian or object modeling found in urban traffic. BARK and CommonRoads support scriptable, and HighwayEnv custom scenarios through code re-implementation, respectively. SMARTS allows to model urban traffic with traffic lights and traffic signs. However, as the focus is not on graphical realism, those scenarios can not reach the visual diversity of traffic scenes in CARLA, which is necessary to train and evaluate vision based AD stacks.
@@
MetaDrive [10] is a training framework that builds on the Panda3D engine. It offers a wide range of features such as procedural road generation and it incorporates scenarios from different traffic datasets. However, scenarios can not be specified by behavioral building blocks or by predefined sub-scenarios. Moreover, its engine limits the structural variety of environments that can be achieved in CARLA.
@@
DI Drive Gym [11] and MACAD Gym [12] are both training frameworks designed around CARLA. DI Drive Gym focuses on single-agent autonomous driving systems and interfaces with ScenarioRunner scenarios, lacking support for training multi-agent systems or scenario generation via sampling. MACAD Gym supports multi-agent training but offers limited scenario generation capabilities, allowing only manual specification of initial conditions.
@@
MATS-Gym in contrast, supports scenario scripting and sampling, realistic sensors, physics, and rendering, urban as well as highway traffic, and (multi-agent) reinforcement learning. Our comparison focuses on relevant aspects for learning-based approaches in realistic urban traffic scenarios. We acknowledge that our criteria may not be complete, and there might be aspects that could be relevant in other use cases. However, we tried to cover a broad spectrum of frameworks that could be used for training and evaluation.
@@
Curriculum Learning. Generating safety-critical scenarios is an active area of research and has been addressed in numerous works in recent years [13]. However, training agents in complex environments requires presenting scenarios at the right difficulty, matching their capabilities to generate meaningful learning signals. Curriculum Learning (CL) and Auto CL (ACL) methods have emerged to progressively expose agents to more complex environments. While traditional CL involves manually designing curricula, ACL automates this process by dynamically presenting suitable scenarios based on agent performance. For an overview, we refer to [14].
@@
Previous works also applied ACL to AD scenarios [15, 16, 17]. In this work, we adopt Unsupervised Environment Design for AD, a recent variation of the ACL methodology where task creation itself is treated as an optimization problem, allowing for more flexible scenario adaptions.
III Background
We model the environment as a Partially Observable Stochastic Game (POSG), defined as a tuple:
$\mathcal{G}=\langle\mathcal{N},\mathcal{S},\{\mathcal{A}_{i}\}_{i\in\mathcal{N}},\{\mathcal{O}_{i}\}_{i\in\mathcal{N}},\mathcal{T},\mathcal{R},\mathcal{Z},\gamma\rangle.$
It consists of the set of agents $\mathcal{N}=\{1,2,\dots,N\}$ and the set of environment states $\mathcal{S}$. For each agent $i\in\mathcal{N}$, $\mathcal{A}_{i}$ denotes the set of actions, and $\mathcal{O}_{i}$ is the set of observations. The joint action space is defined as $\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$, the joint observation space is $\mathcal{O}=\mathcal{O}_{1}\times\mathcal{O}_{2}\times\cdots\times\mathcal{O}_{N}$. The state dynamics are given by $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$, where $\mathcal{T}(s^{\prime}|s,\mathbf{a})$ represents the probability of transitioning from state $s$ to state $s^{\prime}$ given the joint action $\mathbf{a}\in\mathcal{A}$. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}$ specifies the rewards for each agent, where $\mathcal{R}(s,\mathbf{a})=(R_{1}(s,\mathbf{a}),R_{2}(s,\mathbf{a}),\dots,R_{N}(s,\mathbf{a}))$. The observation function $\mathcal{Z}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{O})$ gives the probability $\mathcal{Z}(\mathbf{o}|s,\mathbf{a})$ of receiving joint observation $\mathbf{o}\in\mathcal{O}$ in state $s$ following joint action $\mathbf{a}$. The discount factor $\gamma\in[0,1]$ represents the relative importance of future rewards. At each time step, the system is in a state $s\in\mathcal{S}$, which is not directly observable. Each agent selects an action $a_{i}\in\mathcal{A}_{i}$ based on their observation $o_{i}\in\mathcal{O}_{i}$ by following a stochastic policy $\pi_{i}:\mathcal{O}_{i}\mapsto\Delta(\mathcal{A}_{i})$. The joint action $\mathbf{a}=(a_{1},a_{2},\dots,a_{N})$ is applied, resulting in a new state $s^{\prime}\in\mathcal{S}$ according to the transition function $\mathcal{T}$, and each agent receives a reward ${R_{i}}(s,\mathbf{a})$ and a new observation according to $\mathcal{Z}$. The value function for a policy $\pi$ is the expected sum of discounted future rewards:
$V_{i}^{\pi}(s)=\mathbb{E}_{\mathcal{T}}\left[\sum_{t=0}^{H}\gamma^{t}R_{i}(s_{t},\mathbf{a}_{t})\mid s_{0}=s,a^{i}\sim\pi_{i}\right].$
The goal for agent $i$ is to find a policy $\pi$ that maximizes $V_{i}^{\pi}(s)$ over the horizon $H$. For simplicity, we will drop the agent indices in the remainder of the paper.
We allow that the POSG can be modified at runtime by parametrizing the environment dynamics and the set of agents. This can be formalized as a Underspecified POSG (analogous to [18]):
$\mathcal{G}^{\Theta}=\langle\mathcal{N}^{\Theta},\mathcal{S},\{\mathcal{A}_{i}\}_{i\in\mathcal{N}},\{\mathcal{O}_{i}\}_{i\in\mathcal{N}},\mathcal{T}^{\Theta},\mathcal{R},\mathcal{Z},\gamma,\Theta\rangle,$
where $\Theta$ is a set of parameters that can be chosen by the scenario generator $\tilde{\pi}\in\Delta(\Theta)$. Unsupervised Environment Design (UED) algorithms seek to find generators that maximize the regret of a policy $\pi$:
$\max_{\tilde{\pi}}\mathbb{E}_{\theta\sim\tilde{\pi}}\big{[}\max_{\pi^{\prime}}V^{\pi^{\prime}}-V^{\pi}\big{]},$
@@ III Background
the difference of the maximum achievable return and the actual return. This yields a suitable objective for curriculum design: low regret scenarios indicate that the policy already performs well, while high regret entails difficult but solvable scenarios. Recent works have explored different algorithms for UED. Prioritized Level Replay (PLR) utilizes domain randomization (DR) to generate parameters and a replay buffer to resample levels proportional to their regret, a measure of suboptimality [19]. PAIRED [18] optimizes the level generator via reinforcement learning (RL) to maximize regret, while REPAIRED [4] combines PLR with the level generation approach from PAIRED in a Dual Curriculum Design (DCD) framework. Other approaches use evolutionary algorithms to optimize the generator and include a level editor for environment modifications [20, 21]. In contrast, MATS-Gym takes advantage of the Scenic’s parametrizable scenario specifications, as a way to build an adaptive scenario generation procedure, and adopts a dual-curriculum design algorithm similar to REPAIRED [4].
IV MATS-Gym
@@
MATS-Gym is a multi-agent training and evaluation framework that allows to generate diverse traffic scenarios in CARLA, a high-fidelity traffic simulator. The framework is designed to reconcile scenario execution engines, such as Scenic [2] and ScenarioRunner [22], with multi-agent training frameworks. It offers comprehensive infrastructure for agent-state retrieval, providing ground truth information such as position, velocity vectors, and key traffic events like violations and collisions. MATS-Gym supports a versatile infrastructure for configuring sensor-suites for agents, akin to the CARLA AD Challenge, which supports complex sensors, such as cameras, LiDAR, and Radar. MATS-Gym extends the CARLA BirdEyeView [23] framework for bird’s-eye view observations, by constructing occupancy gridmaps around agents encoding lane markings, traffic lights, signs, and other road users. The agents can also access road-network information and vector-based representations of the map.
We account for different types of tasks, by providing action spaces at various levels of abstraction. On the lowest level, agents can issue throttle, breaking and steering commands at high frequencies. On a higher level, agents can instead provide a target waypoint which is tracked by a PID controller for smooth driving trajectories. Finally, we also provide a discrete action space for high level driving commands, suitable for route planning, behavior prediction, or multi-agent interaction. Furthermore, MATS-Gym allows for defining custom tasks and offers pre-defined tasks related to autonomous driving, including route following, infraction avoidance, and driving comfort, crucial for designing reward functions in reinforcement learning approaches.
IV-A Scenario Based Curriculum Generation
@@
In the following, we will outline our approach to use Scenic’s parametrizable scenario specifications as a way to build an adaptive scenario generation procedure. We adopt a dual-curriculum design algorithm based on REPAIRED [4], a recent algorithm combining PLR with an adaptive scenario generator. Algorithm 1 outlines this approach. At each iteration, we either generate a new scenario by sampling environment parameters from generator $\tilde{\pi}$ or we sample a replay scenario from the PLR buffer. The buffer samples scenarios with a probability proportional to their regret and a staleness metric, as described in the original approach. Then, a trajectory is sampled from the environment to update the policy $\pi_{\phi}$ parametrized by $\phi$ (e.g. a neural network) using any RL algorithm (e.g. PPO [24]). In the next step, we compute the estimated regret for the current scenario to update the buffer and the environment generator. Differently than the original, we do not use an antagonist agent to estimate the regret. Instead, we use the Maximum Monte Carlo [4] regret estimator which approximates the regret to avoid the expensive inner optimization routine:
$\textsc{Regret}(\pi)=\max_{\pi^{\prime}}V^{\pi^{\prime}}-V^{\pi}\approx\frac{1}{H}\sum_{t=0}^{H}R_{\text{max}}-V^{\pi}(s_{t}).$ (1)
This regret formulation compares the maximum achieved return for a scenario with the expectation of the returns for the current policy. Furthermore, for the scenario generator optimization, we use an off-the shelf black-box optimization algorithm [25].
@@
1:  Input: policy $\pi_{\phi}$, scenario generator $\tilde{\pi}$, buffer $\Lambda$
@@
5:        Generate scenario parameters: $\theta\sim\tilde{\pi}$
@@
8:        Sample replay scenario: $\theta\sim\Lambda$
@@
10:     Sample batch $\tau$ with $\theta$: $\{(o_{i},a_{i},r_{i},o^{\prime}_{i})\}^{N}_{i=0}\sim\mathcal{T}_{\pi_{\phi}}^{\theta}$
11:     Policy gradient step: $\phi\leftarrow\phi+\alpha\,\nabla_{\phi}J^{\pi_{\phi}}_{\textsc{PPO-Loss}}(\tau)$
12:     Estimate regret $\hat{R}\approx\textsc{Regret}(\pi)$ using Eq. 1
13:     Update scenario buffer: $\Lambda\leftarrow(\theta,\hat{R})$
14:     if scenario was generated then
15:        $\tilde{\pi}\leftarrow\textsc{Black-Box-Opt}(\theta,\hat{R})$